OpenStudy (anonymous):

Can someone please help me with Stats, I would like to know why this answer is the right one? *Which is relative better a score of 68 on a psychology test or a score of 47 on an economics test? scores on the psychology test have a mean of 86 and a standard deviation of 6. scores on the economics test have a mean of 64 and a standard deviation of 8. I choose: The Economic test score is relatively better because its Z score is greater than the Z score for the psychology test score. Please help!! Thanks!

OpenStudy (anonymous):

@jim_thompson5910, Hi!! I know this is a silly question, but do i have to solve it first? Please help me!

jimthompson5910 (jim_thompson5910):

Z score for psych test z = (x - mu)/sigma z = (68 - 86)/6 z = -3 ------------------------ Z score for econ test z = (x - mu)/sigma z = (47 - 64)/8 z = -2.125 So you are correct, the econ z score is larger, which means it's relatively better.

OpenStudy (anonymous):

OOh nice, Thank you!!! You are my life saver!!...jijij!!

jimthompson5910 (jim_thompson5910):

yw